# Hierarchical Transformer Preconditioning for Interactive Physics Simulation

### 13 May 2026 v1

@@ 6.2. Spectrally Biased Probe Vectors
An isotropic Gaussian probe places equal expected energy on every eigenmode of MAMA (Sec. 3), so the resulting gradient signal is also white in the probe’s eigenbasis. The blocks of our preconditioner, however, are not all at the same spatial scale: leaf-diagonal blocks resolve fine, high-frequency structure over $L$ adjacent nodes, while an off-diagonal tile of size SLSL resolves much lower-frequency, larger-scale structure. A spectrally white probe distribution therefore distributes the gradient signal unevenly across these block scales — the high-frequency components, which the fine-scale (diagonal) blocks are tuned to, dominate the signal; the coarse-scale (off-diagonal, large-$S$) blocks receive proportionally weaker gradients. The downstream effect is that blocks at different scales saturate at different times during training, with the coarse blocks plateauing late and limiting overall convergence.
@@ 6.3. Training Setup
We train with AdamW under a reduce-on-plateau schedule and global gradient clipping. Training contexts (graph, AA, masks, padded sizes, smoothed probes) are precomputed once and cached on disk; at each step a mini-batch is drawn at random and padded to a common node count. We set the number of probe vectors to $K_{z}=\max(64,\lceil\sqrt{N}\rceil)$, balancing gradient noise against compute. The model is compiled with torch.compile. Preconditioner weights and the apply path use float32; only the PCG scalar accumulators (dot products, residual norms, step sizes) are computed in float64, which we found is sufficient to prevent residual drift on stiff systems without paying for full mixed-precision GEMMs.
@@ 7.2. Main Performance
Fig. 6 traces how convergence actually looks on a challenging frame — a closed cross-shaped barrier with stiff density contrast chosen to maximize long-range coupling — across unpreconditioned CG, Jacobi, IC, AMG, and ours, with the right-hand side supported only on thin density interfaces. By k=1k\!=\!1, Jacobi and IC have damped the residual only locally around $\mathrm{supp}(b)$, while AMG and ours have already attenuated it across the whole domain — the visual signature of multiscale transport that the highway buffers implement (§4.3). AMG matches the spread but pays in V-cycle synchronization per iteration; ours runs the inner loop as a single CUDA-Graph capture. Iteration counts on this frame ($288/1097/334/19/384$ for unprec/Jacobi/IC/AMG/ours) span nearly two orders of magnitude, but per-iteration cost reverses the ranking for AMG and IC. Standard deviations across the 20 test frames track condition number rather than NN: Jacobi and unprec. CG show $\sigma/\mu\!\approx\!50$$100\%$ at $N\!\geq\!8\,192$ across the $[5,100]$ contrast range, while ours stays at $\leq\!21\%$ even at N=16 384N\!=\!16\,384.
@@ 7.3. Training dynamics and ablations
Fig. 4 tracks $\mathcal{L}_{\cos}$, the SAI loss on the same checkpoints, and PCG iteration count across training. The three curves move together until step $\sim\!8\,000$, after which the SAI loss rises from $\sim\!10^{-3}$ to $\sim\!0.5$ while $\mathcal{L}_{\cos}$ keeps falling in lockstep with PCG iterations — direct evidence of Prop. 6.1: the model is moving the eigenvalues of MAMA away from $\|A\|$ in pointwise terms while tightening the relative cluster wherever angular alignment is easiest. Fig. 2 confirms this on the spectrum at N=1 024N\!=\!1\,024: SAI delivers a $16\times$ $\kappa$ reduction with the cluster anchored near $\|A\|$, the same architecture trained with $\mathcal{L}_{\cos}$ delivers $68\times$ with the cluster wherever it pleased — a $4.3\times$ gap attributable to the loss alone. Total wall-clock training is $16.2$ min on a single H200 (24 30024\,300 steps); the model overtakes GPU Jacobi by step $\sim\!2\,000$ ($\sim\!1.3$ min) and drops below $200$ PCG iterations by step $\sim\!12\,000$ (Fig. 7 corroborates the link from probe-space alignment to spectral clustering).
@@
Width $d$, depth $n_{l}$, and highway connections each move total solve time non-trivially. Default (d=128d\!=\!128, $n_{l}\!=\!3$, hw on, $\sim\!2.6$M parameters): (i) shrinking depth $n_{l}\!=\!3\!\to\!1$ cuts inference $2.6\times$ but doubles total solve time because iterations rise $2.5\times$ (additional layers are needed to compose information routed through the highway tokens); (ii) removing highways raises PCG iterations $2.3\times$ at N=2 048N\!=\!2\,048 and the penalty grows with NN; (iii) d=64d\!=\!64 is $28\%$ slower overall, d=256d\!=\!256 is competitive at small NN but raises mean PCG iterations from $105$ to $191$ when averaged over $N\!\in\!\{2048,4096,8192\}$. Table 5 summarizes the core sweep results; full rows remain in Supplementary, Table LABEL:SX-supp-tab:ablation.
@@ 8. Discussion and Future Work
The recipe — a weak-admissibility $\mathcal{H}$-matrix prior, a scale-invariant cosine-Hutchinson objective, and a single-graph apply path — is most useful where it targets: stiff, every-frame-different SPD systems with a hard real-time budget. The architecture itself depends on nothing fluid-specific, only a loose spatial ordering of degrees of freedom, and extends directly to 3D (Fig. 1). We expect the largest gains to persist on other SPD families with (i) geometric locality, (ii) frame-to-frame coefficient changes, and (iii) hard real-time budgets (implicit viscosity/diffusion, soft-body and contact dynamics) and smaller gains where a single matrix is reused long enough for heavy classical setup to amortize. Two known limitations are worth flagging: (a) training pre-sizes the partition to a maximum NN, so pushing past it currently requires retraining — a dynamic-partition variant (constant leaf count at $O(N\log N)$ rather than constant leaf size at $\Theta(N)$, §4.2) removes this ceiling at the cost of one extra pooling pass per layer; (b) the $\mathcal{H}$-matrix prior assumes some spatial locality of AA under its indexing, and degrades on operators without a natural spatial coordinate (power-grid Laplacians, social-network matrices) or where the far-field rank does not decay with separation.

### 14 May 2026 v2

@@ 6.2. Spectrally Biased Probe Vectors
An isotropic Gaussian probe places equal expected energy on every eigenmode of $MA$ (Sec. 3), so the resulting gradient signal is also white in the probe’s eigenbasis. The blocks of our preconditioner, however, are not all at the same spatial scale: leaf-diagonal blocks resolve fine, high-frequency structure over $L$ adjacent nodes, while an off-diagonal tile of size $SL$ resolves much lower-frequency, larger-scale structure. A spectrally white probe distribution therefore distributes the gradient signal unevenly across these block scales — the high-frequency components, which the fine-scale (diagonal) blocks are tuned to, dominate the signal; the coarse-scale (off-diagonal, large-$S$) blocks receive proportionally weaker gradients. The downstream effect is that blocks at different scales saturate at different times during training, with the coarse blocks plateauing late and limiting overall convergence.
@@ 6.3. Training Setup
We train with AdamW under a reduce-on-plateau schedule and global gradient clipping. Training contexts (graph, $A$, masks, padded sizes, smoothed probes) are precomputed once and cached on disk; at each step a mini-batch is drawn at random and padded to a common node count. We set the number of probe vectors to $K_{z}=\max(64,\lceil\sqrt{N}\rceil)$, balancing gradient noise against compute. The model is compiled with torch.compile. Preconditioner weights and the apply path use float32; only the PCG scalar accumulators (dot products, residual norms, step sizes) are computed in float64, which we found is sufficient to prevent residual drift on stiff systems without paying for full mixed-precision GEMMs.
@@ 7.2. Main Performance
Fig. 6 traces how convergence actually looks on a challenging frame — a closed cross-shaped barrier with stiff density contrast chosen to maximize long-range coupling — across unpreconditioned CG, Jacobi, IC, AMG, and ours, with the right-hand side supported only on thin density interfaces. By k=1k\!=\!1, Jacobi and IC have damped the residual only locally around $\mathrm{supp}(b)$, while AMG and ours have already attenuated it across the whole domain — the visual signature of multiscale transport that the highway buffers implement (§4.3). AMG matches the spread but pays in V-cycle synchronization per iteration; ours runs the inner loop as a single CUDA-Graph capture. Iteration counts on this frame ($288/1097/334/19/384$ for unprec/Jacobi/IC/AMG/ours) span nearly two orders of magnitude, but per-iteration cost reverses the ranking for AMG and IC. Standard deviations across the 20 test frames track condition number rather than $N$: Jacobi and unprec. CG show $\sigma/\mu\!\approx\!50$$100\%$ at $N\!\geq\!8\,192$ across the $[5,100]$ contrast range, while ours stays at $\leq\!21\%$ even at N=16 384N\!=\!16\,384.
@@ 7.3. Training dynamics and ablations
Fig. 4 tracks $\mathcal{L}_{\cos}$, the SAI loss on the same checkpoints, and PCG iteration count across training. The three curves move together until step $\sim\!8\,000$, after which the SAI loss rises from $\sim\!10^{-3}$ to $\sim\!0.5$ while $\mathcal{L}_{\cos}$ keeps falling in lockstep with PCG iterations — direct evidence of Prop. 6.1: the model is moving the eigenvalues of $MA$ away from $\|A\|$ in pointwise terms while tightening the relative cluster wherever angular alignment is easiest. Fig. 2 confirms this on the spectrum at N=1 024N\!=\!1\,024: SAI delivers a $16\times$ $\kappa$ reduction with the cluster anchored near $\|A\|$, the same architecture trained with $\mathcal{L}_{\cos}$ delivers $68\times$ with the cluster wherever it pleased — a $4.3\times$ gap attributable to the loss alone. Total wall-clock training is $16.2$ min on a single H200 (24 30024\,300 steps); the model overtakes GPU Jacobi by step $\sim\!2\,000$ ($\sim\!1.3$ min) and drops below $200$ PCG iterations by step $\sim\!12\,000$ (Fig. 7 corroborates the link from probe-space alignment to spectral clustering).
@@
Width $d$, depth $n_{l}$, and highway connections each move total solve time non-trivially. Default (d=128d\!=\!128, $n_{l}\!=\!3$, hw on, $\sim\!2.6$M parameters): (i) shrinking depth $n_{l}\!=\!3\!\to\!1$ cuts inference $2.6\times$ but doubles total solve time because iterations rise $2.5\times$ (additional layers are needed to compose information routed through the highway tokens); (ii) removing highways raises PCG iterations $2.3\times$ at N=2 048N\!=\!2\,048 and the penalty grows with $N$; (iii) d=64d\!=\!64 is $28\%$ slower overall, d=256d\!=\!256 is competitive at small $N$ but raises mean PCG iterations from $105$ to $191$ when averaged over $N\!\in\!\{2048,4096,8192\}$. Table 5 summarizes the core sweep results; full rows remain in Supplementary, Table LABEL:SX-supp-tab:ablation.
@@ 8. Discussion and Future Work
The recipe — a weak-admissibility $\mathcal{H}$-matrix prior, a scale-invariant cosine-Hutchinson objective, and a single-graph apply path — is most useful where it targets: stiff, every-frame-different SPD systems with a hard real-time budget. The architecture itself depends on nothing fluid-specific, only a loose spatial ordering of degrees of freedom, and extends directly to 3D (Fig. 1). We expect the largest gains to persist on other SPD families with (i) geometric locality, (ii) frame-to-frame coefficient changes, and (iii) hard real-time budgets (implicit viscosity/diffusion, soft-body and contact dynamics) and smaller gains where a single matrix is reused long enough for heavy classical setup to amortize. Two known limitations are worth flagging: (a) training pre-sizes the partition to a maximum $N$, so pushing past it currently requires retraining — a dynamic-partition variant (constant leaf count at $O(N\log N)$ rather than constant leaf size at $\Theta(N)$, §4.2) removes this ceiling at the cost of one extra pooling pass per layer; (b) the $\mathcal{H}$-matrix prior assumes some spatial locality of $A$ under its indexing, and degrades on operators without a natural spatial coordinate (power-grid Laplacians, social-network matrices) or where the far-field rank does not decay with separation.